OpenStudy (anonymous):

In 1995 the life expectancy of a male in a certain country was 63.6 years In 1999 it was 67.4 years Let E represent the life expectancy in year t and let t represent the number of years since 1995 The linear function E(t) that fits the data is E(t)= t + (round to the nearest tenth) Use the function to predict the life expectancy of males in 2006 E(11)= (round to the nearest tenth)

OpenStudy (anonymous):

i love predicting past events

OpenStudy (anonymous):

are you sure it says \[E(t)=t+...\]

OpenStudy (anonymous):

just think of finding the equation for the line through the points \[(0,63.6),(4,67.4)\]

OpenStudy (anonymous):

e(t)=

OpenStudy (anonymous):

the slope of the line is \[\frac{67.4-63.6}{4}=\frac{3.8}{4}=.95\]

OpenStudy (anonymous):

which is another way of saying the linear model says in 4 years it increased by 3.8 so average in crease is .95 per year

OpenStudy (anonymous):

so \[E(t)=.95t+63.6\] since you started counting at 63.6

OpenStudy (anonymous):

2006 is 11 years after 1995 so find \[E(11)=.95\times 11+63.6\]

OpenStudy (anonymous):

can you round that to the nearest 10th please?